Circular No. 6816

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

1SAX J0104.5+5828 R. Ricci and M. Smith, BeppoSAX Science Operation Center, Tele-

spazio, Rome; P. Soffitta and L. Piro, Istituto Astrofisica Spaziale, CNR, Rome; P. Giommi and S. Rebecchi, BeppoSAX Science Data Center (SDC); J. M. Muller, Space Research Organization of the Netherlands, Utrecht, and SDC; and G. Gennaro, BeppoSAX Operational Control Center, report on behalf of the BeppoSAX team: “On Jan. 28.039583 UT, BeppoSAX WFC no. 2 detected an x-ray flare from a previously unknown source (now designated 1SAX J0104.5+5828). The event lasted ∼ 100 s with approximately equal rise and decay time, and it had a peak flux of 0.2 Crab in the range 2–28 keV. No signal is present between 40 and 700 keV. The pulse profile appears to be not structured, with a flat-topped peak. The position is α = 1^{h}04^{m}35^{s}, δ = +58^{o}28′.8 (equinox 2000.0) — i.e., at a galactic latitude of ∼ −4^{o}. Due to the non-optimal satellite attitude configuration at the moment of the burst, the derived error radius is 10′. No catalogued x-ray sources are present in the error circle.”

COMET 55P/TEMPEL-TUTTLE L. Jorda, Max Planck Institute for Aeronomie; J. Lecacheux, Meudon

Observatory; F. Colas, Bureau des Longitudes, Paris; E. Frappa, Planetarium de Saint-Etienne; and P. Laques, Bagneres-de-Bigorre, France, report that images of the nuclear region of comet 55P taken in the past few days show a jet of matter whose brighter inner section (within 4′′ of the nucleus) moves with an amplitude of 25^{o}, yielding possible rotation periods of 15.31 ± 0.03 and 14.79 ± 0.02 hr. Some selected images are presented

at http://www.bdl.fr/s2p/comete/tt/tt e.html.

Visual m_{1} and coma-diameter estimates (B = binoculars): Jan. 15.04 UT, 7.4, 25′ (R. Keen, Mt. Thorodin, CO, 7×50 B); 17.72, 7.5, 19′ (M. Lehký, Hradec Králové, Czech Rep., 10×80 B); 21.05, 7.4, ∼ 14′ (J. Bortle, Stormville, NY, 10×50 B); 25.3, 7.4, 15′ (S. J. O’Meara, Volcano, HI, 10×50 B); 30.08, 7.9, 12′ (G. W. Kronk, Troy, IL, 20×80 B).

COMET 29P/SCHWASSMANN-WACHMANN 1 This comet increased in brightness by > 3 mag between Jan. 14.25

and 24.25 UT in CCD V frames obtained by R. Casas and J. Licandro at the Observatorio del Teide, Tenerife. Total visual magnitude estimates with 0.41-m reflectors: Jan. 25.47, 12.4 (A. Hale, Cloudcroft, NM; nearly stellar); 26.48, 12.5 (Hale); 29.82, 12.7 (A. Pearce, Nedlands, W. Australia).

1998 January 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green